# COALA: Numerically Stable and Efficient Framework for Context-Aware Low-Rank Approximation

### v2

@@ 3 Weighted Low-Rank Approximation
where $S=(XX^{\top})^{1/2}$ is the unique positive definite square root of $XX^{\top}$, $U\Sigma V^{\top}$ is the SVD of $WS$, and $\Sigma_{r}$ is the matrix obtained from $\Sigma$ by setting the last nrn-r singular values to zero.
@@ 4 Inversion-Free Solution
is attained at $W^{\prime}=U_{r}U_{r}^{\top}W$, where $U_{r}$ consists of the first $r$ left singular vectors of the matrix $WX$.
@@
It is well-known that the minimizer of $\|A-B\|_{F}$ under the constraint $\operatorname{rank}(B)\leq r$ is given by $B=U_{r}U_{r}^{\top}A$, where $U_{r}$ contains the first $k$ left singular vectors of $A$ (see Corollary 2 for details). Substituting back for $A$ and $B$, we have $B=U_{r}U_{r}^{\top}A=U_{r}U_{r}^{\top}WX$, implying
@@ 5 Weighted Low-Rank Approximation with Regularization
3Form the matrix $X^{\prime}=[X\quad\sqrt{\mu}\,I]$, where $I$ is the $n\times n$ identity matrix.
@@
Let $W\in\mathbb{R}^{m\times n}$ and $X\in\mathbb{R}^{n\times k}$. Suppose that $X$ has $\operatorname{rank}(X)=k\geq r$ and that the singular values of $WX$ satisfy $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, where $\sigma_{i}(\cdot)$ denotes the $i$-th largest singular value. Let $W_{0}=U_{r}U_{r}^{\top}W$ denote the solution to the problem (3), and let $W_{\mu}$ denote the solution to the regularized problem (4). then the following estimate holds:
@@ 6.1 Model compression
where $\lambda$ serves as a hyperparameter controlling the adjustment. This step is crucial because different layers of large language models exhibit substantially different norms of the weight matrices $W$, calibration matrices $X$, and their products $WX$, see, e.g., [19]. Moreover, we compare these two strategies, analyzing how the metric depends on adaptive and non-adaptive choices of $\mu$ across layers, see Figure 4. The Mistral‑7B‑Instruct model was selected due to its pronounced variation of layer‑wise norms.
@@ 60. Guidelines:
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
@@ Appendix B SVD-LLM Method
5Compute the singular value decomposition of $WS$:
@@ Appendix C Basics of Low-Rank Approximation
by keeping only the top $r$ singular values $\sigma_{1},\dots,\sigma_{r}$ in $\Sigma$, along with the corresponding columns of $U$ and $V$. Then $A_{r}$ is a best rank-$r$ approximation to $A$ in the Frobenius norm, i.e.
@@
can be obtained using the formula $A_{k}=U_{k}U_{k}^{\top}A$ or $A_{k}=AV_{k}V_{k}^{\top}$, where the SVD of matrix $A$ is given by
@@
Let $A\in\mathbb{R}^{m\times n}$ be a matrix such that its $r$-th and $(r+1)$-th singular values satisfy $\sigma_{r}(A)\neq\sigma_{r+1}(A)$. Let $E\in\mathbb{R}^{m\times n}$ be a perturbation matrix, and define $\hat{A}=A+E$. Let $U_{r}\in\mathbb{R}^{m\times r}$ and $\hat{U}_{r}\in\mathbb{R}^{m\times r}$ be matrices whose columns consist of the first $r$ left singular vectors of $A$ and $\hat{A}$, respectively. Then,
@@ Appendix D Convergence Proofs for the Full-Rank Regularization Problem
Let $W\in\mathbb{R}^{m\times n}$ and $X\in\mathbb{R}^{n\times k}$. Suppose that $X$ has full row rank (i.e., $\operatorname{rank}(X)=n$) and that the singular values of $WX$ satisfy $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, where $\sigma_{i}(\cdot)$ denotes the $i$-th largest singular value. Then, the solution $W_{0}$ to the problem (1) is unique. Furthermore, if $W_{\mu}$ denotes the solution to the regularized problem (4) , then the following estimate holds:
@@
The uniqueness of $W_{0}$ follows from the fact that if $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, then the rank-$r$ low-rank approximation $Y_{r}$ of the matrix $WX$ is unique (by Eckart-Young-Mirsky Theorem 3). Hence, $W_{0}$ is a solution if and only if $W_{0}X=Y_{r}$. Moreover, since the kernel of $X$ is empty, if such a matrix $W_{0}$ exists, it must be unique. However, by Proposition 1 , such a matrix indeed exists.
@@ Appendix E Convergence Proofs (Without the Full-Rank Condition)
where $U_{0}$ denotes the first $r$ left singular vectors of $WX$, and $H_{0}=(XX^{\top})^{1/2}$. Analogously, using Proposition 3 , we have
@@
Here we use the fact that the left singular vectors of a matrix $A$ coincide with the eigenvectors vectors (same that left singular in this case) of $AA^{\top}$.
@@ Appendix G Examples
We fixed all dimensional parameters, left and right singular vectors of the matrix $WX$, as well as all singular values except for the $r$-th and $(r+1)$-th ones. Then, we varied the difference between $\sigma_{r}(WX)$ and $\sigma_{r+1}(WX)$. The convergence rate of the regularized solution to the unregularized solution with respect to this gap is presented Figure 6. We observe that the dependence on the gap is intrinsic to the problem and that we catch the correct asymptotic behaviour in our theoretical bound in the full rank case.

### v3

@@ 3 Weighted Low-Rank Approximation
where $S=(XX^{\top})^{1/2}$ is the unique positive definite square root of $XX^{\top}$, $U\Sigma V^{\top}$ is the SVD of WSWS, and $\Sigma_{r}$ is the matrix obtained from $\Sigma$ by setting the last nrn-r singular values to zero.
@@ 4 Inversion-Free Solution
is attained at $W^{\prime}=U_{r}U_{r}^{\top}W$, where $U_{r}$ consists of the first $r$ left singular vectors of the matrix WXWX.
@@
It is well-known that the minimizer of $\|A-B\|_{F}$ under the constraint $\operatorname{rank}(B)\leq r$ is given by $B=U_{r}U_{r}^{\top}A$, where $U_{r}$ contains the first $k$ left singular vectors of AA (see Corollary 2 for details). Substituting back for AA and $B$, we have $B=U_{r}U_{r}^{\top}A=U_{r}U_{r}^{\top}WX$, implying
@@ 5 Weighted Low-Rank Approximation with Regularization
3Form the matrix $X^{\prime}=[X\quad\sqrt{\mu}\,I]$, where II is the $n\times n$ identity matrix.
@@
Let $W\in\mathbb{R}^{m\times n}$ and $X\in\mathbb{R}^{n\times k}$. Suppose that the singular values of WXWX satisfy $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, where $\sigma_{i}(\cdot)$ denotes the $i$-th largest singular value. Let $W_{0}=U_{r}U_{r}^{\top}W$ denote the solution to the problem (3), and let $W_{\mu}$ denote the solution to the regularized problem (4). then the following estimate holds:
@@ 6.1 Model compression
where $\lambda$ serves as a hyperparameter controlling the adjustment. This step is crucial because different layers of large language models exhibit substantially different norms of the weight matrices $W$, calibration matrices $X$, and their products WXWX, see, e.g., [19]. Moreover, we compare these two strategies, analyzing how the metric depends on adaptive and non-adaptive choices of $\mu$ across layers, see Figure 4. The Mistral‑7B‑Instruct model was selected due to its pronounced variation of layer‑wise norms.
@@ 60. Guidelines:
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
@@ Appendix B SVD-LLM Method
5Compute the singular value decomposition of WSWS:
@@ Appendix C Basics of Low-Rank Approximation
by keeping only the top $r$ singular values $\sigma_{1},\dots,\sigma_{r}$ in $\Sigma$, along with the corresponding columns of $U$ and $V$. Then $A_{r}$ is a best rank-$r$ approximation to AA in the Frobenius norm, i.e.
@@
can be obtained using the formula $A_{k}=U_{k}U_{k}^{\top}A$ or $A_{k}=AV_{k}V_{k}^{\top}$, where the SVD of matrix AA is given by
@@
Let $A\in\mathbb{R}^{m\times n}$ be a matrix such that its $r$-th and $(r+1)$-th singular values satisfy $\sigma_{r}(A)\neq\sigma_{r+1}(A)$. Let $E\in\mathbb{R}^{m\times n}$ be a perturbation matrix, and define $\hat{A}=A+E$. Let $U_{r}\in\mathbb{R}^{m\times r}$ and $\hat{U}_{r}\in\mathbb{R}^{m\times r}$ be matrices whose columns consist of the first $r$ left singular vectors of AA and $\hat{A}$, respectively. Then,
@@ Appendix D Convergence Proofs for the Full-Rank Regularization Problem
Let $W\in\mathbb{R}^{m\times n}$ and $X\in\mathbb{R}^{n\times k}$. Suppose that $X$ has full row rank (i.e., $\operatorname{rank}(X)=n$) and that the singular values of WXWX satisfy $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, where $\sigma_{i}(\cdot)$ denotes the $i$-th largest singular value. Then, the solution $W_{0}$ to the problem (1) is unique. Furthermore, if $W_{\mu}$ denotes the solution to the regularized problem (4) , then the following estimate holds:
@@
The uniqueness of $W_{0}$ follows from the fact that if $\sigma_{r}(WX)\neq\sigma_{r+1}(WX)$, then the rank-$r$ low-rank approximation $Y_{r}$ of the matrix WXWX is unique (by Eckart-Young-Mirsky Theorem 3). Hence, $W_{0}$ is a solution if and only if $W_{0}X=Y_{r}$. Moreover, since the kernel of $X$ is empty, if such a matrix $W_{0}$ exists, it must be unique. However, by Proposition 1 , such a matrix indeed exists.
@@ Appendix E Convergence Proofs (Without the Full-Rank Condition)
where $U_{0}$ denotes the first $r$ left singular vectors of WXWX, and $H_{0}=(XX^{\top})^{1/2}$. Analogously, using Proposition 3 , we have
@@
Here we use the fact that the left singular vectors of a matrix AA coincide with the eigenvectors vectors (same that left singular in this case) of $AA^{\top}$.
@@ Appendix G Examples
We fixed all dimensional parameters, left and right singular vectors of the matrix WXWX, as well as all singular values except for the $r$-th and $(r+1)$-th ones. Then, we varied the difference between $\sigma_{r}(WX)$ and $\sigma_{r+1}(WX)$. The convergence rate of the regularized solution to the unregularized solution with respect to this gap is presented Figure 6. We observe that the dependence on the gap is intrinsic to the problem and that we catch the correct asymptotic behaviour in our theoretical bound in the full rank case.